OpenStudy (k.binks):

Write the equation of a hyperbola with vertices (-5, 0) and (5, 0) and foci (-7, 0) and (7, 0)

OpenStudy (anonymous):

You know how to form the equation of hyperbola?

OpenStudy (anonymous):

Into this form \[\frac{ x^2 }{ a^2 } - \frac{ y^2 }{ b^2 } =1\]

OpenStudy (k.binks):

Ok I can do that, I'm just getting my formulas mixed up. thanks

OpenStudy (anonymous):

yw